OpenStudy (anonymous):

If 8/9 = x/15, then x=?

OpenStudy (ganpat):

8/9 = x/15, x = (8 * 15) / 9 = 40/3

OpenStudy (anonymous):

\[\frac 39 = \frac{x}{15}\]\[\LARGE Cross \space Multiply!\]\[3(15) = 9x\]\[x = ?\]

OpenStudy (anonymous):

But it is 8/9 not 3/9

OpenStudy (anonymous):

Oops. lol \[\frac 89 = \frac{x}{15}\]\[8(15) = 9x\]

OpenStudy (anonymous):

It ends up as 13 and 1/3. So x= that?

OpenStudy (anonymous):

You should get an improper fraction.

OpenStudy (anonymous):

Alright. Yup. \[13\frac13 \space or \space \frac{40}{3}\]

OpenStudy (anonymous):

THat's what x equals.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

np :)